# Properties of leveled spatial graphs

### v1

@@ Abstract
Leveled spatial graphs are studied regarding other spatial graph properties, including being free or paneled. New graph invariants are defined for an abstract graph permitting a leveled embedding: the level number and the hamiltonian level number for hamiltonian graphs. The level numbers for complete graphs and complete bipartite graphs are obtained.
@@ 1 Introduction
The family of leveled spatial graphs was originally introduced in order to study the possibility of embedding a spatial graph cellular on a surface [1]. In this paper, we investigate the property of a spatial graph to have a leveled embedding. We give an equivalent definition for a leveled embedding to the original one and characterize the abstract graphs with a leveled embedding. We show that all leveled embeddings are free and we compare leveled and paneled (also known as flat) embeddings.
Moreover, we introduce new graph invariants for (hamiltonian) graphs which admit a (hamiltonian) leveled embedding, called the (hamiltonian) level number. These invariants provide a measure on how far a graph is from being planar. We investigate how spatial graphs being leveled compares with other spatial graph properties. In particular, we study the relation between the level number and other graph invariants that minimize the decomposition of a graph in planar subgraphs, namely the thickness [10] and book thickness [3] of a graph.
We characterize graphs with low level number and determine the level number of complete and complete bipartite graphs.
@@ 2 Preliminaries
All graphs are connected undirected simple finite graphs. The fragments (or bridges) of a graph $G$ with respect to a cycle $C$ of $G$ are the closures of the connected components of $G-C$ [9]. Two fragments are conflicting if they either have pairs of endpoints on $C$ that alternate, or have at least three endpoints in common. The conflict graph of a graph $G$ with respect to a cycle $C$ is the graph that has a vertex for each fragment of $G$ with respect to $C$, and two vertices are adjacent if and only if the corresponding fragments conflict. A connected graph is hamiltonian if it contains a cycle that visits every vertex exactly once. Consequently, all fragments of a hamiltonian graph with respect to the hamiltonian cycle are closures of edges or loops. An abstract graph $G$ is planar if it has an embedding on a sphere or a plane. It is maximal planar if $G$ is planar but adding any edge to $G$ makes it non-planar. $G$ is outerplanar if it can be embedded on the sphere such that all vertices lie on the boundary of one region. The chromatic number of a graph $G$ is the minimum number of colors that are required to color every vertex of $G$ such that adjacent vertices have different colors. A graph embedding is an embedding $f:G\rightarrow S^{3}$ of a graph $G$ in $S^{3}$ up to ambient isotopy. The corresponding spatial graph $\mathcal{G}$ is the image of this embedding up to ambient isotopy. A spatial graph is trivial if it embeds in $S^{2}$.
For $n\geq 1$, an $n$-book consists of a line $L$, called the spine of the book, and $n$ distinct half-planes, the pages of the book, with boundary $L$. An $n$-book embedding of a graph $G$ is an embedding $\gamma$ of $G$ in an $n$-book such that each vertex of $G$ lies on the spine $L$ and each edge lies in the interior of one page of the book. The vertices of $G$ occur in some order along the spine of the book and this sequence is called the printing cycle of the embedding [3]. A spatial graph $\mathcal{G}$ is paneled (or flat) if every cycle $C$ of $\mathcal{G}$ bounds an embedded open disk in $S^{3}$ disjoint from $\mathcal{G}$. A spatial graph $\mathcal{G}$ is free if the fundamental group of its complement $\pi_{1}(S^{3}\setminus\mathcal{G})$ is a free group. A diagram of a spatial graph $\mathcal{G}$ is the image of a projection of $\mathcal{G}$ onto $\mathbb{R}^{2}$ such that all intersections in the image are transversal, no vertices and at most two points are mapped to a crossing, and each double point is assigned as an over- or under-crossing. A diagram is reduced if no Reidemeister moves of type 1 and 2 can be applied to reduce its number of crossings. A leveled spatial graph $\mathcal{G}$ is a connected spatial graph that contains an unknotted cycle $C$ called its spine such that each fragment $f$ of $\mathcal{G}$ with respect to $C$ can be embedded in a disk $D_{f}$ whose boundary is identified with $C$, with interior $\mathring{D_{f}}$ disjoint from $\mathcal{G}-f$, in a way that no two disks $D_{f}$ and $D_{g}$ for distinct fragments $f$ and $g$ intersect [1].
@@
In this paper, we use a different definition of leveled spatial graphs (definition 3), which facilitates the study of its relation with other spatial graph properties. We prove the equivalence of the two definitions in proposition 1.
@@
Let $G$ be a graph. We say that an embedding of $G$ is semi-leveled if there exists an unknotted cycle $C$ in $G$, called the spine, and a diagram $D$ of the embedded graph in which no fragment $f$ with respect to $C$ has self-crossings, crossings with $C$, or both over and undercrosses with another fragment. The diagram $D$ is called a semi-leveled diagram of the semi-leveled spatial graph $\mathcal{G}$. We say that $\mathcal{G}$ is semi-leveled with respect to spine $C$ when we want to stress the choice of the spine of the semi-leveled embedding.
@@
Let $\mathcal{G}$ be a semi-leveled spatial graph and let $D$ be a reduced semi-leveled diagram of $\mathcal{G}$, and such that every fragment with respect to the spine $C$ lies in the bounded connected component of $\mathbb{R}^{2}\setminus C$. We say that $D$ is divided in levels if there exists a partition =$\mathscr{L}=\{L_{1},L_{2},\dots,L_{n}\}$ of all the fragments, called the level partition, such that the following holds: $L_{1}$ is the set of fragments that do not cross over any other fragment, and for $i>1$, the set $L_{i}$ consists of fragments that cross over at least one fragment $f\in L_{i-1}$ and only cross over fragments belonging to $L_{k}$, with $k<i$. The fragments belonging to $L_{i}$ are called fragments of level $i$ and $n$ is the number of levels of the embedding of $\mathcal{G}$.
@@
A connected spatial graph $\mathcal{G}$ has a leveled embedding if it has a semi-leveled embedding with a semi-leveled diagram $D$ that is divided in levels. (fig. 1)
@@
Clearly, leveled implies semi-leveled but the opposite does not hold true, i.e. there exist semi-leveled diagrams which cannot be divided in levels. Indeed, the spatial graph $\mathcal{G}$ containing the torus knot $T(3,2)$ shown in fig. 2 is semi-leveled but not leveled. Indeed, inspecting the six hamiltonian cycles of $\mathcal{G}$ as possible spines, we can see that it is either not possible to find a level partition $\mathscr{L}$ or that there is no semi-leveled diagram.
@@
Given a connected spatial graph $\mathcal{G}$, definition 3 is equivalent to the definition of a leveled spatial graph given in the text above Definition 1.
@@
If $\mathcal{G}$ has a semi-leveled embedding with semi-leveled diagram $D$, then every fragment $f$ with respect to the spine $C$ is a planar subgraph of $\mathcal{G}$ without self-crossings or crossings with $C$. Therefore, $f$ can be embedded on a disk $D_{f}$ whose boundary is identified with $C$. Since $D$ is divided in levels, the interior of $D_{f}$ is disjoint from $\mathcal{G}-f$ and any other $D_{g}$, for $g$ another fragment of $\mathcal{G}$. For the other implication, order the disks in a sequence $D_{1},\dots D_{d}$ such that disks $D_{f}$ that together bound a ball in the complement of the graph are labeled consecutively modulo the number of disks $d$. Consider the diagram $D$ given by embedding the spine $C$ on a plane $\mathbb{R}^{2}$ and projecting the disks $D_{f}$ to the bounded connected component of $\mathbb{R}^{2}\setminus C$. Without loss of generality, the diagram can be assumed to be reduced, and such that the fragment of disk $D_{1}$ is not crossing over any other fragment. The diagram $D$ is semi-leveled, because no fragment has self-crossings, crossings with $C$, or both over- and undercrossings with another fragment by construction. Define a level partition in the following way: Label the fragments of $\mathcal{G}$ by the index of the disks they belong to. Assign $f_{1}$ and all fragments that do not cross over any other fragment in $D$ to level 1. Assign to level 2 the fragment with smallest label that conflicts with a fragment assigned to level 1 and does not conflict with fragments assigned to level 2. Continue in the same way until all fragments are assigned to a level. The defined partition is a level partition and therefore $D$ is divided in levels.∎
@@ 3 Characterization and stability of leveled embeddings
Leveled embeddings are not “stable” in the sense that not all cycles of the graph can be chosen as spines, not even in the case that all fragments with respect to the spine are trivial. For example, on the left of fig. 3 a hamiltonian leveled spatial graph is shown. The blue hamiltonian cycle cannot act as a spine since all diagrams in which the blue hamiltonian does not have self-crossings, two of its fragments have both over- and under-crossings.
The problem in fig. 3 is that the spatial graph contains a knot. However, we can still find examples of knotless leveled spatial graphs with a cycle whose all fragments are trivial, such that the cycle cannot act as spine, see fig. 4. It is known that changing the spine of a leveled embedding affects its number of levels [1].
@@ 4 Leveled spatial graphs are free
In proposition 3 we prove that the family of leveled embeddings is contained in the family of free embeddings. To prove this result, we need first to show a sufficient condition for a spatial graph to be free.
@@
The converse is not true. Indeed, consider the trefoil knot with one unknotting tunnel drawn in fig. 5. This is a free spatial graph. Since the contraction of edges does not change the complement of the spatial graph, contracting an edge which is part of the knot results in a free spatial graph that is not trivial. However, the only spanning tree of the obtained graph is its vertex and therefore no spanning tree can be found that could be contracted to obtain a trivial graph.
@@
However, Makino and Suzuki showed that there exists a sequence of edge contractions and vertex splittings between any pair of spatial graphs that have the same neighbourhood [7]. This allows us to state the following:
@@
Given that $\mathcal{G}$ is leveled, we can embed the fragment $f$ of $\mathcal{G}$ in a disk $D_{f}$ such that $\partial D_{f}$ coincides with $C$ and the interior of $D_{f}$ does not intersect $\mathcal{G}$ except in $f$. If the fragment $f$ contains a vertex of $\mathcal{G}$ that is not a vertex of the spine, we can contract edges of $f$ inside $D_{f}$ until there are no vertices of $f$ outside the spine. As the obtained spatial graph $\mathcal{G}^{\prime}$ is hamiltonian leveled, we can again embed each fragment $f^{\prime}$ of $\mathcal{G}^{\prime}$ on a disk $D_{f}^{\prime}$ disjoint from $\mathcal{G}^{\prime}-f^{\prime}$. Therefore, we can assume $\mathcal{G}$ to be hamiltonian leveled without loss of generality. Note that for a hamiltonian leveled graph $\mathcal{G}$, $f$ splits $D_{f}$ into two disks, $D_{f,1}$ and $D_{f,2}$. Moreover, since $C$ is unknotted, there exists a disk $D_{C}$ such that $\partial D_{C}$ coincides with $C$ and the interior of $D_{C}$ does not intersect $\mathcal{G}$. By lemma 1, it is enough to find a spanning tree $T$ such that $\mathcal{G}/T$ is trivial to prove the statement. Let $C$ be the spine of $\mathcal{G}$ and let $e$ be any edge in $C$. Consider $T=C-e$, then $T$ is a spanning tree for $\mathcal{G}$. Contracting $T$, the boundary of $D_{f}$ becomes the edge $e$, while the fragment $f$ becomes a loop based at the only vertex of the bouquet $\mathcal{G}/T$. Only one of $D_{f,1},D_{f,2}$ remains a disk, without loss of generality we can assume it is $D_{f,1}$ for every fragment $f$, while the other becomes a pinched disk. In a similar way, $D_{C}$ remains a disk with boundary given by $e$. As the collection of disks $D_{C},D_{f,1}$ for each fragment $f$ of $\mathcal{G}$ have interiors which are pairwise disjoint, the spatial graph $\mathcal{G}/T$ is paneled. By (2.2) of [8], $\mathcal{G}/T$ is trivial since it is a paneled embedding of a planar abstract graph. Therefore, $\mathcal{G}$ is free. ∎

### v2

@@ Abstract
We investigate the property of a spatial graph of having a leveled embedding and characterize the abstract graphs with this property. We show that all leveled embeddings are free and we compare leveled and paneled (also known as flat) embeddings.
Two new graph invariants are introduced: the level number, an invariant for graphs that admit a leveled embedding, and the Hamiltonian level number, an invariant for Hamiltonian graphs. These invariants provide a measure on how far a graph is from being planar. We study the relation between the (Hamiltonian) level number and other graph invariants that minimize the decomposition of a graph in planar subgraphs, namely the thickness and the book thickness of a graph.
We characterize graphs with low level number and determine both the level number and the Hamiltonian level number of complete graphs and of complete bipartite graphs.
@@ 1 Introduction
The family of leveled spatial graphs was originally introduced in order to study the possibility of embedding a spatial graph cellular on a surface [1]. Intuitively, a leveled spatial graph consists of an unknotted cycle with subgraphs attached to it, arranged so that the subgraphs lie on stacked disks whose common boundary is the cycle. Definitions are given in section 2.
We give a new definition of leveled spatial graph and show that it is equivalent to the previous definition (proposition 1). With the new definition it becomes easier to characterize which abstract graphs, i.e. graphs without a specified embedding, admit a leveled embedding. This is done in proposition 2 and corollary 1 where we show that every hamiltonian graph can be embedded as a leveled spatial graph. Moreover, the new definition helps in comparing leveledness with other important spatial graph properties that relate to algebraic properties of the embedding: freeness and paneledness. In proposition 3 we show that all leveled spatial graphs are free.
It is known that the family of paneled spatial graphs is a subfamily of free graph embeddings, since Robertson, Seymour and Thomas proved that a spatial graph is paneled if and only if all its subgraphs are free ([10]). We compare paneledness and leveldness and show in theorem 2 that if a paneled graph embedding admits a spine, i.e. a cycle such that every fragment is a planar subgraph, then the embedding is leveled. A consequence of this result is corollary 1 which states that all Hamiltonian paneled spatial graphs are leveled.
Leveled spatial graphs can be viewed as a generalization of planar graphs, in the sense that a leveled spatial graph consists of stacked planar subgraphs whose union is the whole graph. Because of this interpretation, we define two new graph invariants: the level number and the Hamiltonian level number. These invariants are defined as the lowest number of levels a spatial graph (respectively, a Hamiltonian graph) admits ranging over all of its possible leveled embeddings (respectively, Hamiltonian leveled embeddings). The (Hamiltonian) level number measures how far a graph is from being planar, hence we compare the new invariants with two other invariants which also minimize the decomposition of a graph in planar subgraphs: the thickness [12] and the book thickness [3] of a graph. Even though level number, thickness and book thickness are similar, we show in proposition 6 and the discussion thereafter that they are all distinct. Finally, we determine the level numbers of complete graphs and of complete bipartite graphs in proposition 7 and theorem 3.
@@ 2 Preliminaries
For the concepts not defined here, we refer the reader to books of topological graph theory and of general topology, for example [6] and [9]. All graphs in this paper are connected undirected simple finite graphs.
The fragments (or bridges) of a graph $G$ with respect to a cycle $C$ of $G$ are the closures of the connected components of GCG-C [11]. Two fragments are conflicting if they either have pairs of endpoints on $C$ that alternate, or have at least three endpoints in common. The conflict graph of a graph $G$ with respect to a cycle $C$ is the graph that has a vertex for each fragment of $G$ with respect to $C$, and two vertices are adjacent if and only if the corresponding fragments conflict.
@@
A connected graph is Hamiltonian if it contains a cycle that visits every vertex exactly once. Consequently, all fragments of a Hamiltonian graph with respect to the Hamiltonian cycle are closures of edges or loops. An abstract graph $G$ is planar if it has an embedding on a sphere or a plane. It is maximal planar if $G$ is planar but adding any edge to $G$ makes it non-planar. $G$ is outerplanar if it can be embedded on the sphere such that all vertices lie on the boundary of one region. The chromatic number of a graph $G$ is the minimum number of colors that are required to color every vertex of $G$ such that adjacent vertices have different colors.
A graph embedding is an embedding $f:G\rightarrow S^{3}$ of a graph $G$ in the three-dimensional sphere $S^{3}$, up to ambient isotopy [^1] For a formal definition of ambient isotopy we refer the reader to [9]. Informally, an ambient isotopy of a spatial graph allows to continuously deform edges and vertices of the graph along with the surrounding space $S^{3}$. In this way, self-intersections of the graph and shrinking knots to a point are avoided.. The corresponding spatial graph $\mathcal{G}$ is the image of this embedding up to ambient isotopy. A spatial graph is trivial if it embeds in the 2-dimensional sphere $S^{2}$.
For $n\geq 1$, an $n$-book consists of a line $L$, called the spine of the book, and $n$ distinct half-planes, the pages of the book, with boundary $L$. An $n$-book embedding of a graph $G$ is an embedding of $G$ in an $n$-book such that each vertex of $G$ lies on the spine $L$ and each edge lies in the interior of one page of the book. The vertices of $G$ occur in some order along the spine of the book and this sequence is called the printing cycle of the embedding [3].
A spatial graph $\mathcal{G}$ is paneled (or flat) if every cycle $C$ of $\mathcal{G}$ bounds an embedded open disk in $S^{3}$ disjoint from $\mathcal{G}$. A spatial graph $\mathcal{G}$ is free if the fundamental group of its complement $\pi_{1}(S^{3}\setminus\mathcal{G})$ is a free group, i.e. a group whose generators do not have any relations.
A diagram of a spatial graph $\mathcal{G}$ is the image of a projection of $\mathcal{G}$ onto the Euclidean plane $\mathbb{R}^{2}$ such that all intersections in the image are transversal, no vertices and at most two points are mapped to a crossing, and each double point is assigned as an over- or under-crossing. A diagram is reduced if no Reidemeister moves of type 1 and 2 can be applied to reduce its number of crossings. These moves are shown in Figure 1. A cycle is unknotted if it can be embedded as the equator of a sphere $S^{2}$.
In our previous work [1] we defined leveled spatial graphs as follows:
Definition 1.
A leveled spatial graph $\mathcal{G}$ is a connected spatial graph that contains an unknotted cycle $C$ called its spine such that each fragment $f$ of $\mathcal{G}$ with respect to $C$ can be embedded in a disk $D_{f}$ whose boundary is identified with $C$, with interior $\mathring{D_{f}}$ disjoint from $\mathcal{G}-f$, in a way that no two disks $D_{f}$ and $D_{g}$ for distinct fragments $f$ and $g$ intersect.
For the special case of the spine to be a Hamiltonian cycle of the graph, we introduce Hamiltonian leveled spatial graphs:
Definition 2.
@@
Figure 1: Reidemeister moves of type 1 (above) and type 2 (below).
In the present paper, we define leveled spatial graphs alternatively (definition 5) as the spatial graphs with properties defined in definition 3 and definition 4. This facilitates the study of leveledness in relation to other spatial graph properties. We prove the equivalence of the two definitions in proposition 1.
@@
Let $G$ be a graph. We say that an embedding of $G$ is semi-leveled if there exists an unknotted cycle $C$ in $G$, called the spine, and a diagram $D$ of the embedded graph in which no fragment $f$ with respect to $C$ has self-crossings, crossings with $C$, or both over- and under-crossings with another fragment. The diagram $D$ is called a semi-leveled diagram of the semi-leveled spatial graph $\mathcal{G}$. We say that $\mathcal{G}$ is semi-leveled with respect to spine $C$ when we want to stress the choice of the spine of the semi-leveled embedding.
@@
Let $\mathcal{G}$ be a semi-leveled spatial graph and let $D$ be a reduced semi-leveled diagram of $\mathcal{G}$, such that every fragment with respect to the spine $C$ lies in the bounded connected component of $\mathbb{R}^{2}\setminus C$. We say that $D$ is divided in levels if there exists a partition $\mathscr{L}=\{L_{1},L_{2},\dots,L_{n}\}$ of all the fragments, called the level partition, such that the following holds: $L_{1}$ is the set of fragments that do not cross over any other fragment, and for $i>1$, the set $L_{i}$ consists of fragments that cross over at least one fragment $f\in L_{i-1}$ and only cross over fragments belonging to $L_{k}$, with $k<i$. The fragments belonging to $L_{i}$ are called fragments of level $i$ and $n$ is the number of levels of the embedding of $\mathcal{G}$.
@@
A connected spatial graph $\mathcal{G}$ has a leveled embedding if it has a semi-leveled embedding with a semi-leveled diagram $D$ that is divided in levels. (Figure 2)
@@
Clearly, a leveled embedding is semi-leveled but the opposite does not hold true, i.e. there exist semi-leveled diagrams which cannot be divided in levels. Indeed, the spatial graph $\mathcal{G}$ containing the torus knot $T(3,2)$ shown in Figure 3 is semi-leveled but not leveled. Indeed, inspecting the six Hamiltonian cycles of $\mathcal{G}$ as possible spines, we can see that it is either not possible to find a level partition $\mathscr{L}$ or that there is no semi-leveled diagram.
@@
Definition 1 and definition 5 are equivalent.
@@
If $\mathcal{G}$ has a semi-leveled embedding with semi-leveled diagram $D$, then every fragment $f$ with respect to the spine $C$ is a planar subgraph of $\mathcal{G}$ without self-crossings or crossings with $C$. Therefore, $f$ can be embedded on a disk $D_{f}$ whose boundary is identified with $C$. Since $D$ is divided in levels, the interior of $D_{f}$ is disjoint from $\mathcal{G}-f$ and any other $D_{g}$, for $g$ another fragment of $\mathcal{G}$. For the other implication, order the disks in a sequence $D_{1},\dots,D_{d}$ such that disks $D_{f}$ that together bound a ball in the complement of the graph are labeled consecutively modulo the number of disks $d$. Consider the diagram $D$ given by embedding the spine $C$ on a plane $\mathbb{R}^{2}$ and projecting the disks $D_{f}$ to the bounded connected component of $\mathbb{R}^{2}\setminus C$. Without loss of generality, the diagram can be assumed to be reduced, and such that the fragment of disk $D_{1}$ is not crossing over any other fragment. The diagram $D$ is semi-leveled, because no fragment has self-crossings, crossings with $C$, or both over- and under-crossings with another fragment by construction. Define a level partition in the following way: Label the fragments of $\mathcal{G}$ by the index of the disks they belong to. Assign $f_{1}$ and all fragments that do not cross over any other fragment in $D$ to level 1. Assign to level 2 the fragment with smallest label that conflicts with a fragment assigned to level 1 and does not conflict with fragments assigned to level 2. Continue in the same way until all fragments are assigned to a level. The defined partition is a level partition and therefore $D$ is divided in levels.∎
@@ 3 Characterization and stability of leveled embeddings
Leveled embeddings are not “stable” in the sense that not all cycles of the graph can be chosen as spines, not even in the case that all fragments with respect to the spine are trivial. For example, on the left of Figure 4 a Hamiltonian leveled spatial graph is shown. The blue Hamiltonian cycle cannot act as a spine since all diagrams in which the blue Hamiltonian does not have self-crossings, two of its fragments have both over- and under-crossings.
The problem in Figure 4 is that the spatial graph contains a knot. However, we can still find examples of knotless leveled spatial graphs with a cycle whose all fragments are trivial, such that the cycle cannot act as spine, see Figure 5. It is known that changing the spine of a leveled embedding affects its number of levels [1].
@@ 4 Leveled spatial graphs are free
In proposition 3 we prove that the family of leveled embeddings is contained in the family of free embeddings. To prove this result, we need first to show a sufficient condition for a spatial graph to be free. A bouquet is a graph consisting of a single vertex. The graph $G/e$, where $e$ is an edge of $G$ with distinct endpoints, is obtained form $G$ by contracting the edge $e$ and letting its endpoints coincide. In a similar fashion, the graph $G/T$, where $T$ is a spanning tree of $G$, is obtained from $G$ contracting each edge of $T$ one at a time.
@@
The converse is not true. Indeed, consider the trefoil knot with one unknotting tunnel drawn in Figure 6. This is a free spatial graph. Since the contraction of edges does not change the complement of the spatial graph, contracting an edge which is part of the knot results in a free spatial graph that is not trivial. However, the only spanning tree of the obtained graph is its vertex and therefore no spanning tree can be found that could be contracted to obtain a trivial graph.
@@
However, Makino and Suzuki showed that there exists a sequence of edge contractions and vertex splittings between any pair of spatial graphs that have the same neighbourhood [8], where the neighbourhood of a spatial graph $\mathcal{G}$ is a tubular neighbourhood of its edges glued to tubular neighbourhoods of its vertices in such a way that $\mathcal{G}$ is embedded in the neighbourhood. This allows us to state the following:
@@
Given that $\mathcal{G}$ is leveled, we can embed the fragment $f$ of $\mathcal{G}$ in a disk $D_{f}$ such that $\partial D_{f}$ coincides with $C$ and the interior of $D_{f}$ does not intersect $\mathcal{G}$ except in $f$. If the fragment $f$ contains a vertex of $\mathcal{G}$ that is not a vertex of the spine, we can contract edges of $f$ inside $D_{f}$ until there are no vertices of $f$ outside the spine. As the obtained spatial graph $\mathcal{G}^{\prime}$ is Hamiltonian leveled, we can again embed each fragment $f^{\prime}$ of $\mathcal{G}^{\prime}$ on a disk $D_{f}^{\prime}$ disjoint from $\mathcal{G}^{\prime}-f^{\prime}$. Therefore, we can assume $\mathcal{G}$ to be Hamiltonian leveled without loss of generality. Note that for a Hamiltonian leveled graph $\mathcal{G}$, $f$ splits $D_{f}$ into two disks, $D_{f,1}$ and $D_{f,2}$. Moreover, since $C$ is unknotted, there exists a disk $D_{C}$ such that $\partial D_{C}$ coincides with $C$ and the interior of $D_{C}$ does not intersect $\mathcal{G}$. By lemma 1, it is enough to find a spanning tree $T$ such that $\mathcal{G}/T$ is trivial to prove the statement. Let $C$ be the spine of $\mathcal{G}$ and let $e$ be any edge in $C$. Consider T=CeT=C-e, then $T$ is a spanning tree for $\mathcal{G}$. Contracting $T$, the boundary of $D_{f}$ becomes the edge $e$, while the fragment $f$ becomes a loop based at the only vertex of the bouquet $\mathcal{G}/T$. Only one of $D_{f,1},D_{f,2}$ remains a disk, without loss of generality we can assume it is $D_{f,1}$ for every fragment $f$, while the other becomes a pinched disk. In a similar way, $D_{C}$ remains a disk with boundary given by $e$. As the collection of disks $D_{C},D_{f,1}$ for each fragment $f$ of $\mathcal{G}$ have interiors which are pairwise disjoint, the spatial graph $\mathcal{G}/T$ is paneled. By (2.2) of [10], $\mathcal{G}/T$ is trivial since it is a paneled embedding of a planar abstract graph. Therefore, $\mathcal{G}$ is free. ∎